Could someone check my answer for this one-

The formula for measuring drag force on a car is Fd=1/2pv^2ACd where Fd is the drag force measured in Newtons, p is density, v is velocity, A is area and Cd is the drag coefficient. Assuming that all the factors are measured in SI units, what units should Cd have to make the formula dimensionally homogeneous?^
Fd=kgm/s^2
p=kg/m^3
v=m/s
A=m^2
Cd=?
Therefore kg*m/s^2 / 1/2*kg/m^3*m/s*m^2
Cancelling the kg and m^2 and m^3
Seem to leave m / s * 1/2
Does this mean Cd has no units or could I inverse 1/2 and 1/s to make m^s2

Answer by Alan3354(69443)

The formula for measuring drag force on a car is Fd=1/2pv^2ACd where Fd is the drag force measured in Newtons, p is density, v is velocity, A is area and Cd is the drag coefficient. Assuming that all the factors are measured in SI units, what units should Cd have to make the formula dimensionally homogeneous?^
Fd=kgm/s^2
p=kg/m^3
v=m/s
A=m^2
Cd=?
Therefore kg*m/s^2 / (1/2*kg/m^3*m/s*m^2) *** It's v^2
Cd = 2 no units
-----------------
Cancelling the kg and m^2 and m^3
Seem to leave m / s * 1/2
Does this mean Cd has no units or could I inverse 1/2 and 1/s to make m^s2